A particle of mass m describes a circle of radius r. The centripetal acceleration of particlr is 4/r^2 . Msgnitude of momentum(p) = ?

Grade:8

2 Answers

Maharshi Sen
6 years ago
2m/[(r)^0.5] towards center of circle.
[(v^2)]/r = 4/[r^2]
v = 2/[r^0.5]
This is in accordance with centripetal accelaration.  angular accelaration = 0. hopes it helps you.
Anjali
5 years ago
Acceleration of particle(centripetal) = 4/r^2
              v^2/r = 4/r^2
                     v = 2/√r
Now , momentum = m×v
                                = m×2/√r
                                = 2m/√r